An automotive manufacturer wants to know the proportion of new car buyers who prefer foreign cars over domestic.

Step 1 of 2:

Suppose a sample of 1453 new car buyers is drawn. Of those sampled, 363 preferred foreign over domestic cars. Using the data, estimate the proportion of new car buyers who prefer foreign cars. Enter your answer as a fraction or a decimal number rounded to three decimal places.
Answer #1

estimated value of the proportion of new car buyers who prefer foreign cars = 363 / 1453 = 0.250 (Ans)

**If this answers do not match or any kind of confusion you have please comment